While preparing for breakfast, your mother throws the egg vertically upward creating a speed of 8.0 m/s at a height of 2 m. The mass of the egg is 1.5 g. Find its mechanical energy.


1
Expert's answer
2021-03-17T16:52:58-0400

"E=KE+PE=\frac{mv^2}{2}+mgh="


"=\frac{0.0015\cdot 8^2}{2}+0.0015\cdot9.8\cdot2=0.0774\ (J)=77.4\ (mJ)" . Answer